OpenStudy (anonymous):

Evaluate 15÷3·(-2) + |-10|

OpenStudy (anonymous):

0

OpenStudy (anonymous):

-10+|(-10)|

OpenStudy (anonymous):

zero

OpenStudy (anonymous):

Order of operations. Multi and Div, left to right, followed by addition and subtraction left to right. 15/3 = 5, 5*(-2) = -10, -10 + |-10| = 0